OpenStudy (anonymous):

Which of the following statements is correct about the two systems of equations?

OpenStudy (anonymous):

how do i do this?

OpenStudy (anonymous):

@mathmate @mathmath333 @mathman101 @HelpOfTheGods

OpenStudy (anonymous):

so definitely x=2 and y=1 in both systems. Now, check by way of working on a piece of paper, you will realise that the first statement is actually false. -10x+19y+1is NOT =[ (2x+y-5)+ 2(-4x+6y+2)= 2x-8x+y+12y-5+4=-6x+13y-1] so you have to test statement by statement, to see if what they are saying is true. Above, I simply said the first eqt in B is NOT EQUAL TO the first equation in A plus 2 times the second equation in A. If you do it all well, if not mistaken, the second statement is TRUE

OpenStudy (tkhunny):

Don't tag multiple people. Show your work, instead.

OpenStudy (anonymous):

ha ha true